# Modeling Diffusion and Permeation Across the Stratum Corneum Lipid Barrier

### v1

@@ 3.3 Challenge 2: Determination of PACF Correlation Times
The challenges associated with extrapolating $D(s)$ to zero frequency were recognized by Hummer, who proposed an alternative approach based on determining the PACF correlation time (eq. 5).28 For purely diffusive dynamics, this method is formally equivalent to the VACF approach, but it is considerably simpler to apply in practice.
In Figure 4, we show selected diffusivities, $D(t)$, as a function of the integration cutoff time used in eq. 5. The PACF method performs well when the solute resides in the aqueous phase, where $D(t)$ reaches a clear plateau between approximately $8\text{\,}\mathrm{ps}$ and $10\text{\,}\mathrm{ps}$. Although noise leads to some fluctuations, it does not significantly alter the predicted diffusivity. In contrast, at the membrane center, the decay of the PACF is substantially slower, and no clear plateau in $D(t)$ is observed even after $10\text{\,}\mathrm{ps}$. These observations are consistent with those reported by Rowley et al. for a DPPC membrane.22
@@ 3.4 Propagator Analysis
Figure 5: Free-energy profiles for the translocation of all solutes across the model stratum corneum membrane, obtained from umbrella sampling (US) and well-tempered metadynamics (WTM).19 The two methods produce consistent overall features, with only minor quantitative differences. Figure adapted with permission from ref. 19.
@@
Having obtained two qualitatively distinct diffusivity profiles from the same set of simulations, it is natural to ask which of the two, if either, is correct. To address this question, we computed the propagators of both diffusion models in the critical membrane-center region, $p(z,t\,|\,z_{0}=0,t_{0}=0)$—that is, the probability of finding a particle initially located at $z_{0}=0$ at position $z$ after a lag time $t$. The corresponding FE profiles were taken from our previous study,19 as shown in Figure 5. For the present analysis, we used the umbrella-sampling (US)–based free-energy profiles. A comparison between the model-derived propagators and those obtained directly from molecular dynamics simulations is presented in Figure 6 for two representative lag times.
At short lag times ($t=$2\text{\,}\mathrm{ps}$\$), the VACF-based model shows quantitative agreement with the simulation data over six orders of magnitude, whereas the PACF approach fails to do so. This finding is consistent with the results reported by Rowley et al.,22 who performed a simplified analysis by approximating the propagator as a normal distribution—an assumption that holds only for lag times long enough to reproduce diffusive dynamics yet short enough that the effects of inhomogeneity remain negligible.
However, it is not obvious that the lag time of $2\text{\,}\mathrm{ps}$ chosen here (or the $5\text{\,}\mathrm{ps}$ used by Rowley et al.) is sufficiently long for the diffusive approximation to hold. Indeed, the failure of the PACF to decay within such short timescales may also be interpreted as a breakdown of this assumption. Subdiffusive behavior has, in fact, been reported for lag times up to at least $64\text{\,}\mathrm{ps}$ for methanol permeation across a POPC membrane.31
@@ 3.5 Impact on Permeabilities
Within the inhomogeneous solubility–diffusion (ISD) model,39, 40 the local resistance of a membrane to permeation, $R(z)$, is expressed in terms of the diffusivity and FE profiles as
@@
We have also shown that the FE profiles obtained from US and WTM differ slightly (Figure 5), even after extensive sampling, although these differences remain within the limits of statistical uncertainty.19 With the diffusivity profiles now available, we can propagate the different FE profiles to evaluate their impact on the resulting transport coefficients. We find good agreement between the results obtained from both methods for water and 6-MHO; for acetone, however, noticeable deviations are observed. Nevertheless, the permeability coefficients for all solutes differ by less than an order of magnitude, indicating that statistical variations between FE methods likely have little to no impact on kinetic model predictions.
More striking are the differences in permeabilities when comparing VACF- and PACF-derived diffusivities. Because the VACF-derived diffusivities are consistently higher than those obtained from the PACF method, the same trend is reflected in the corresponding permeabilities. In absolute terms, deviations reach approximately one order of magnitude. Whether such differences are significant for kinetic models is highly problem specific. In many cases, these discrepancies are likely acceptable, although the final assessment should be guided by a dedicated sensitivity analysis.17
@@ 4 Conclusions
Direct comparison of permeabilities computed from our MD simulations to experimental values (in vitro41, 42 or in vivo43) is not straightforward because our model represents a simplified SC membrane composed of an equimolar mixture of ceramides, free fatty acids, and cholesterol. In contrast, native human SC exhibits substantial lipid heterogeneity, encompassing a broad spectrum of ceramide subclasses and chain-length distributions.44 Nevertheless, the water permeability obtained using PACF-derived diffusivities is within the same order of magnitude as the experimental estimate of Gupta et al. ($7.08\pm 1.18\times 10^{-5}$ $\mathrm{cm}\text{\,}{\mathrm{s}}^{-1}$ ),35 with residual differences most likely reflecting variations between the GROMOS and CHARMM force fields.
@@
Permeability coefficients are often estimated empirically using quantitative structure–activity relationship (QSAR) and continuum diffusion models,45, 46, 47, 48, 49 which correlate measured permeabilities with molecular descriptors such as molar mass, partition coefficients, and bulk diffusion constants. Although such models are powerful for interpolation within their training domains, they do not explicitly capture the molecular-scale interactions and dynamic heterogeneity that govern transport across the SC lipid matrix—especially when applied to molecules that fall outside their original parameterization range. In contrast, MD simulations capture these effects directly by resolving how solutes interact with and move through the heterogeneous lipid environment. For example, despite its larger molar mass relative to water and acetone, 6-MHO exhibits the highest permeability because of its comparatively low free-energy barriers to permeation (Figure 5), underscoring the dominant role of thermodynamic driving forces—rather than molecular size—in determining SC transport.
@@
With respect to indoor air chemistry, several kinetic models—such as KM-BL, KM-FILM, and KM-SUB-Skin-Clothing17—have been developed to describe the dynamics of skin-oil oxidation products across different interfaces. However, these models critically depend on physicochemical parameters that are difficult or impossible to measure experimentally, underscoring the value of molecular-scale simulations like those presented here. For example, Lakey et al. demonstrated that the performance of their KM-FILM model is highly sensitive to the assumed diffusion constants of surface films,50 illustrating how improved microscopic input can directly enhance macroscopic predictive accuracy.
@@
The supporting information includes an assessment of the influence of the harmonic force constant on diffusivity profiles; details on the calculation of bulk diffusivities from mean-squared displacements; the complete set of propagator analyses across different lag times; density profiles used to define membrane boundaries; additional local resistance profiles obtained from all combinations of free-energy and diffusivity methods; and a table summarizing all computed permeability coefficients.

### v2

@@ 3.3 Challenge 2: Determination of PACF Correlation Times
The challenges associated with extrapolating $D(s)$ to zero frequency were recognized by Hummer, who proposed an alternative approach based on determining the PACF correlation time (eq. 5).hummer_position-dependent_2005 For purely diffusive dynamics, this method is formally equivalent to the VACF approach, but it is considerably simpler to apply in practice.
In Figure 4, we show selected diffusivities, $D(t)$, as a function of the integration cutoff time used in eq. 5. The PACF method performs well when the solute resides in the aqueous phase, where $D(t)$ reaches a clear plateau between approximately $8\text{\,}\mathrm{ps}$ and $10\text{\,}\mathrm{ps}$. Although noise leads to some fluctuations, it does not significantly alter the predicted diffusivity. In contrast, at the membrane center, the decay of the PACF is substantially slower, and no clear plateau in $D(t)$ is observed even after $10\text{\,}\mathrm{ps}$. These observations are consistent with those reported by Rowley et al. for a DPPC membrane.gaalswyk_generalized_2016
@@ 3.4 Propagator Analysis
Figure 5: Free-energy profiles for the translocation of all solutes across the model stratum corneum membrane, obtained from umbrella sampling (US) and well-tempered metadynamics (WTM).thomas_insights_2025 The two methods produce consistent overall features, with only minor quantitative differences. Figure adapted with permission from ref. thomas_insights_2025.
@@
Having obtained two qualitatively distinct diffusivity profiles from the same set of simulations, it is natural to ask which of the two, if either, is correct. To address this question, we computed the propagators of both diffusion models in the critical membrane-center region, $p(z,t\,|\,z_{0}=0,t_{0}=0)$—that is, the probability of finding a particle initially located at $z_{0}=0$ at position $z$ after a lag time $t$. The corresponding FE profiles were taken from our previous study,thomas_insights_2025 as shown in Figure 5. For the present analysis, we used the umbrella-sampling (US)–based free-energy profiles. A comparison between the model-derived propagators and those obtained directly from molecular dynamics simulations is presented in Figure 6 for two representative lag times.
At short lag times ($t=$2\text{\,}\mathrm{ps}$\$), the VACF-based model shows quantitative agreement with the simulation data over six orders of magnitude, whereas the PACF approach fails to do so. This finding is consistent with the results reported by Rowley et al.,gaalswyk_generalized_2016 who performed a simplified analysis by approximating the propagator as a normal distribution—an assumption that holds only for lag times long enough to reproduce diffusive dynamics yet short enough that the effects of inhomogeneity remain negligible.
However, it is not obvious that the lag time of $2\text{\,}\mathrm{ps}$ chosen here (or the $5\text{\,}\mathrm{ps}$ used by Rowley et al.) is sufficiently long for the diffusive approximation to hold. Indeed, the failure of the PACF to decay within such short timescales may also be interpreted as a breakdown of this assumption. Subdiffusive behavior has, in fact, been reported for lag times up to at least $64\text{\,}\mathrm{ps}$ for methanol permeation across a POPC membrane.Chipot.Comer.2016.SubdiffusionMembranePermeation
@@ 3.5 Impact on Permeabilities
Within the inhomogeneous solubility–diffusion (ISD) model,Diamond.Katz.1974.InterpretationNonelectrolytePartition, Marrink.Berendsen.1994.SimulationWaterTransport the local resistance of a membrane to permeation, $R(z)$, is expressed in terms of the diffusivity and FE profiles as
@@
We have also shown that the FE profiles obtained from US and WTM differ slightly (Figure 5), even after extensive sampling, although these differences remain within the limits of statistical uncertainty.thomas_insights_2025 With the diffusivity profiles now available, we can propagate the different FE profiles to evaluate their impact on the resulting transport coefficients. We find good agreement between the results obtained from both methods for water and 6-MHO; for acetone, however, noticeable deviations are observed. Nevertheless, the permeability coefficients for all solutes differ by less than an order of magnitude, indicating that statistical variations between FE methods likely have little to no impact on kinetic model predictions.
More striking are the differences in permeabilities when comparing VACF- and PACF-derived diffusivities. Because the VACF-derived diffusivities are consistently higher than those obtained from the PACF method, the same trend is reflected in the corresponding permeabilities. In absolute terms, deviations reach approximately one order of magnitude. Whether such differences are significant for kinetic models is highly problem specific. In many cases, these discrepancies are likely acceptable, although the final assessment should be guided by a dedicated sensitivity analysis.Lakey.Shiraiwa.2019.ImpactClothingOzone
It is important to note here that diffusivity, free energy, and permeation reflect different aspects of transport. A high free-energy barrier (as in the case of water at the bilayer center) reduces the probability of occupancy in that region but does not necessarily imply fast diffusion once a molecule is there. Permeability depends on both thermodynamics and kinetics through eq. 7, and intuition based solely on an unfavourable free-energy environment risks conflating the role of $\Delta F(z)$ with that of $D(z)$. Diffusivity profiles $D(z)$ probe the mobility of a solute conditioned on being at position $z$, not the likelihood of being there in the first place. Within this interpretation, a lower diffusivity $D(z)$ at the membrane center does not mean that water prefers this region, only that its motion is more hindered when present. A surprising consequence of this interplay between thermodynamic and kinetic effects is the relatively high permeability of 6-MHO, despite the fact that its diffusivity profile is generally the lowest of the three solutes.
@@
We conclude this section by noting that a direct molecular interpretation of the resistance profiles in Fig. 8 in terms of the membrane structure in Fig. 1 is non-trivial. The local resistance $R(z)$ reflects the combined effect of free-energy barriers and diffusivity (eq. 7), meaning that structural features do not map one-to-one onto the profile but instead contribute through a convolution of thermodynamic and kinetic influences. Nevertheless, we can conjecture about some qualitative structure–transport relations. The elevated resistance near the headgroup region is consistent with the dense packing and strong hydrogen-bonding environment at lipid–water interfaces, which restricts mobility particularly for polar permeants. In the membrane center, the loosened lipid organization and resulting elevated free volume give rise to conditions that may accelerate local diffusion and contribute to the central resistance dip.
Table 1: Water permeability coefficients of stratum corneum (SC) lipid membranes, obtained from molecular dynamics simulations and experimental compilations. Simulation results are reported together with the corresponding force fields, temperatures, and analysis methods used to compute permeability, which differ across studies and contribute to variability in the reported values. Units are $\mathrm{cm}\text{\,}{\mathrm{s}}^{-1}$.
Reference Permeability Approach Details
Gupta et al. gupta_molecular_2016
$7.08\pm 1.18\times 10^{-5}$ FACF 310 K, GROMOS/SPC
Das et al. das_water_2009
$8.2\times 10^{-8}$ FACF 350 K, OPLS/SPC
Reuter et al. reuter_presence_2025
$1.88\times 10^{-5}$ FACF 303.15 K, CHARMM36/TIP3P
Piasentin et al. piasentin_evaluation_2021
$2.79\pm 0.52\times 10^{-5}$ FACF 303.15 K, CHARMM36/TIP3P
$2.10\pm 0.39\times 10^{-5}$ US + FACF 303.15 K, CHARMM36/TIP3P
$1.89\pm 0.34\times 10^{-5}$ US + VACF 303.15 K, CHARMM36/TIP3P
This work
$5.65\times 10^{-5}$ US + PACF 305.15 K, CHARMM36/TIP3P
$5.48\times 10^{-5}$ WTM + PACF 305.15 K, CHARMM36/TIP3P
$4.15\times 10^{-4}$ US + VACF 305.15 K, CHARMM36/TIP3P
$4.00\times 10^{-4}$ WTM + VACF 305.15 K, CHARMM36/TIP3P
Katritzky et al. katritzky_skin_2006a
$2.95\times 10^{-7}$ Experimental
Chen et al. chen_modeling_2010a
$4.79\times 10^{-7}$ Experimental
  • a
    The reported permeability values are compiled from multiple secondary sources and do not represent primary experimental measurements.
@@ 4 Conclusions
To place these results in context, we compare the water permeability coefficients obtained in this work with values reported in previous studies. The permeabilities derived from PACF-based diffusivities are of the same order of magnitude as those reported by Gupta et al., Piasentin et al., and Reuter et al. gupta_molecular_2016, piasentin_evaluation_2021, Reuter.Lunter.2025.PresenceDifferentCeramide, namely on the order of $10^{-5}$ $\mathrm{cm}\text{\,}{\mathrm{s}}^{-1}$ (see Table 1). Residual discrepancies are most likely attributable to differences in simulation temperature and force-field parametrization. Additional methodological factors—such as system size, thermostat and barostat settings, integration parameters, and overall simulation protocols—may also contribute to variations in the reported permeability coefficients. Importantly, this level of agreement should not be interpreted as a validation of the PACF-based diffusivity profiles. The limitations of the approach discussed here are methodological in nature and are therefore expected to be largely independent of specific computational details.
Direct comparison of permeabilities computed from our MD simulations to experimental values (in vitroanderson_heterogeneity_1988, anderson_solute_1989 or in vivokalia_homogeneous_1996) is not straightforward because our model represents a simplified SC membrane containing only the short periodicity phase (SPP) of the lipid matrix. We deliberately omitted the more structurally complex long periodicity phase (LPP) to first address methodological challenges that arise even in comparatively simple SC models. Improving the realism of atomistic representations is important, but equally essential is ensuring that these methodological limitations are understood and managed if future permeability estimates are to be more robust. We therefore emphasize the need to quantify uncertainty—whether through error bars, bounding estimates, or credibility intervals—and to evaluate how these propagate into higher-scale models.
@@
Permeability coefficients are often estimated empirically using quantitative structure–activity relationship (QSAR) and continuum diffusion models,wilschut_estimating_1995, lian_evaluation_2008, Mitragotri.Roberts.2011, Frasch.Barbero.2013, Wang.Naegel.2023 which correlate measured permeabilities with molecular descriptors such as molar mass, partition coefficients, and bulk diffusion constants. Although such models are powerful for interpolation within their training domains, they do not explicitly capture the molecular-scale interactions and dynamic heterogeneity that govern transport across the SC lipid matrix—especially when applied to molecules that fall outside their original parameterization range. In contrast, MD simulations capture these effects directly by resolving how solutes interact with and move through the heterogeneous lipid environment. For example, despite its larger molar mass relative to water and acetone, 6-MHO exhibits the highest permeability because of its comparatively low free-energy barriers to permeation (Figure 5), underscoring the dominant role of thermodynamic driving forces—rather than molecular size—in determining SC transport.
@@
With respect to indoor air chemistry, several kinetic models—such as KM-BL, KM-FILM, and KM-SUB-Skin-ClothingLakey.Shiraiwa.2019.ImpactClothingOzone—have been developed to describe the dynamics of skin-oil oxidation products across different interfaces. However, these models critically depend on physicochemical parameters that are difficult or impossible to measure experimentally, underscoring the value of molecular-scale simulations like those presented here. For example, Lakey et al. demonstrated that the performance of their KM-FILM model is highly sensitive to the assumed diffusion constants of surface films,lakey_kinetic_2021 illustrating how improved microscopic input can directly enhance macroscopic predictive accuracy.
@@
[…]